OpenStudy (anonymous):

x + 2y = 3 3x - 2y = 5 .What is the solution to the system of equations? A) (4, 3.5) B) (2, 1/2) C) (2, -1) D) (1/2, 2)

Parth (parthkohli):

Add both equations. Solve by elimination.

OpenStudy (anonymous):

Add the two equations to get \(4x=8\), solve for x and substitute to get y.

Parth (parthkohli):

You'll eliminate y, then solve for x. Substitute the value of x and solve for y.

OpenStudy (anonymous):

x+2y=3 3x-y=5 Add both the equations x+3x+2y-2y=3+5 2y cancels out 4x=8 x=4/8 x=2 Now put x =2 in first equation that is x+2y=3 2+2y=3 2y= 3-2 2y=1 y=1/2 therefore your answers are x=2 and y=1/2